OpenStudy (anonymous):

What is the vertex of the graph of y = |x – 6| – 1? What is the vertex of the graph of y = |x – 6| – 1? @Mathematics

OpenStudy (anonymous):

-1,6

jimthompson5910 (jim_thompson5910):

The vertex of y = a|x-h|+k is (h,k) In the case of y = |x – 6| – 1, h = 6 and k = -1 So the vertex is (6, -1)

OpenStudy (anonymous):

6=x -1=y

OpenStudy (anonymous):